Hi, my name is Grace Rogers and I'm having some trouble solving this equation:
|1-4x| +8 =13
I got one solution to be -1, but I can't get the other.
thanks for the help!

Found 2 solutions by richwmiller, ewatrrr:
Answer by richwmiller(17219)

x=3/2 x=-1

+(1-4x) +8 =13
+(1-4x) =5
-4x=4
x=-4/4
x=-1
-(1-4x) +8 =13
-(1-4x) =5
4x-1=5
4x=6
x=6/4
x=3/2

Answer by ewatrrr(24785)

 
Hi
|1-4x| +8 =13
|1-4x| +8 =5
1-4x = 5, x = -1
0r ***
1-4x = -5, x = 6/4 = 3/2